# Embedded Hyperspectral Band Selection with Adaptive Optimization for Image Semantic Segmentation

### v1

@@ Abstract
Hyperspectral band selection plays a pivotal role in remote sensing and image analysis, aiming to identify the most informative spectral bands while minimizing computational overhead. In this paper, we introduce a pioneering approach for hyperspectral band selection that offers an embedded solution, making it well-suited for resource-constrained or real-time applications. Our proposed method, embedded Hyperspectral Band Selection (EHBS), excels in selecting the best bands without the need for prior processing, seamlessly integrating with the downstream task model. This is achieved through the adaptation of the Stochastic Gates (STG) algorithm, originally designed for feature selection, for hyperspectral band selection in the context of image semantic segmentation and the integration of a dynamic optimizer, DoG, which removes the need for the required tuning the learning rate.
To assess the performance of our method, we introduce a novel metric for evaluating band selection methods across different target numbers of selected bands quantified by the Area Under the Curve (AUC). We conduct experiments on two distinct semantic-segmentation hyperspectral benchmark datasets, demonstrating its superiority in terms of its resulting accuracy and its ease of use compared to many common and state-of-the-art methods. Furthermore, our contributions extend beyond the realm of hyperspectral band selection. The adaptability of our approach to other tasks, especially those involving grouped features, opens up promising avenues for broader applications within the realm of deep learning, such as feature selection for feature groups. The demonstrated success on the tested datasets and the potential for application to a variety of tasks underscore the value of our method as a substantial addition to the field of computer vision. It offers new possibilities for feature selection and optimization in complex data analysis scenarios, making it a significant asset for researchers and practitioners alike.
@@ 1 Introduction
While HSI offers notable advantages, it is not without its challenges. One significant drawback lies in the substantial costs associated with the physical sensor hardware. Additionally, HSI incurs increased expenses related to storing, transferring, and analyzing the considerably larger image data generated by HSI. Furthermore, the rise of very large and deep networks for vision has further increased the complexity of models and, consequently, computing costs. Hence, there is a pressing need for algorithms and methods for band selection, i.e., identifying a subset of hyperspectral bands that retains essential information for downstream tasks. Though existing research has explored band selection, the various methods proposed typically involved an unsupervised pre-processing step that is independent of the downstream HSI task and thus the choice of band selection may be sub-optimal.
This paper introduces Embedded Hyperspectral Band Selection (EHBS), a plug-and-play embedded method. EHBS effectively selects the optimal bands without requiring preliminary processing, seamlessly integrating with the downstream task model. EHBS utilizes an existing feature selection algorithm based on stochastic gates that was adopted to the setting of band selection of hyperspectral data. In this study, we showcase the capability of EHBS to dynamically learn optimal bands seamlessly as an integral part of the Convolutional Neural Network (CNN) implementation. This is done in the context of semantic segmentation, a visual task that predicts semantic categories for each pixel in an input image, that has received growing interest in the context of hyperspectral imagery, particularly with the application of deep learning methods. Our results demonstrate that our method outperforms existing band-selection methods on two different hyperspectral semantic segmentation datasets achieving similar accuracy values of the full hyperspectral data by using only around 25% of the bands. Importantly, our approach stands out by being easily applicable to deep-learning tasks over HSI datasets.
@@ 2.1 Hyperspectral Imaging
Hyperspectral imaging captures rich spectral data per pixel, unlike standard imaging with only three spectral samples (Red, Green, Blue). Contemporary hyperspectral systems offer hundreds of spectral bands spanning both visible and invisible spectra, effectively creating a three-dimensional cube composed of two-dimensional grayscale images. This detailed HSI data provides insights beyond the capabilities of regular RGB imaging that can be exploited by deep learning models.
@@
Semantic segmentation is a fundamental HSI task to predict the semantic categories for each pixel of an hyperspectral input image [25]. Semantic segmentation tasks based on aerial and satellite images play an important role in a wide range of applications [20]. In recent years, the successful application of deep learning (DL) in the field of computer vision (CV) has led to a surge of work applying DL methods for data semantic segmentation, resulting in notable achievements and significant advancements.
@@
Deep learning applications for HSI became popular and prevailed in recent years. For example, Zhang et. al [37] used a 3D CNN with transfer learning for aero image segmentation while [29] explored unsupervised hyperspectral unmixing using autoencoders to classify hyperspectral images into 3 labels (Tree,Water,Rock).
@@ 2.2 Feature Selection
Feature selection (FS) methods identify the essential features needed by a machine learning system to perform a downstream task and can be categorized broadly as filter, wrapper, and embedded methods. Filter methods eliminate irrelevant features before model learning, using statistical relevance scores [6, 10]. Wrapper methods determine feature relevance based on model performance [23, 4], but their drawback lies in computational expense [30]. Embedded methods address this by simultaneously learning the model and selecting relevant features within a self-contained and single-train organism. For example, the widely recognized Least Absolute Shrinkage and Selection Operator (LASSO) [34] is an embeded FS algorithm that minimizes loss with an $l_{1}$ constraint but is limited to linear functions. Attempts to extend LASSO using neural networks face challenges with suboptimal gradient descent on $l_{1}$ regularized objectives [15].
STG [36] propose an embedded feature selection neural network scheme. The STG procedure is based on probabilistic relaxation of the $\ell_{0}$ norm of features, or the count of the number of selected features. The $\ell_{0}$-based regularization relies on a continuous relaxation of the Bernoulli distribution; such relaxation allows the STG model to learn the parameters of the approximate Bernoulli distributions via gradient descent. The STG framework simultaneously learns either a nonlinear regression or classification function while selecting a small subset of features and provides an information-theoretic justification for incorporating Bernoulli distribution into feature selection. Figure 1 illustrates the STG model in which the stochastic gates are attached to the $x_{i}$ input features, where the trainable parameter $\mu_{i}$ and a noise component $\epsilon_{i}$ control the choice of the gate being active or not ($z_{i}$).
@@ 2.3 Hyperspectral Band Selection
Band selection (BS) methods are categorized into two main groups. Unsupervised methods operate without using any annotations, relying on information criteria for internal information and dissimilarity from other bands. While unsupervised methods are suitable for unlabeled datasets, their lack of consideration for specific hyperspectral imaging (HSI) task performance may result in sub-optimal band choices. On the other hand, Supervised methods utilize annotated data in the selection process, tailoring band selection to the specific task and considering prediction model performance. This category further divides into before-train (finding the best bands before training the downstream model) and embedded in train (training the band selector as part of the downstream model, as in our method). Following we will describe six different methods that can be classified into five different families: Search, Ranking, Clustering, Sparsity and DL-based. These methods were implemented and used as baseline models to be compared to our proposed model.
@@
Clustering-based methods in band selection involve grouping bands and selecting one representative from each group to minimize redundancy. Often used in conjunction with ranking, these methods aim to choose the highest-ranked representatives from clustered bands [35]. WaLuDI and WaLuMI [27] employ a distance metric based on information theory measures to assess similarity between bands. Their approach minimizes in-cluster variance while maximizing between-cluster variance. The dissimilarity measure calculates the distance between bands using probability functions based on gray-scale pixel values. For instance, the KL divergence is employed to quantify differences between probability distributions, enhancing the effectiveness of band clustering and selection strategies.
@@
Another more recent line of work include DL-based methods aimed to model the nonlinear interdependencies between the various spectral bands. BS-Nets [9] is an embedded unsupervised method that applies a DNN to reconstruct the full HSI image from partially available bands. The networks attention mechanism was used from both spatial and spectral views to infer the best bands combination for reconstruction.
@@ 3.1 Problem statement
The goal of embedded band selection methods is to simultaneously select and Indicator vector $I$ and a model $f_{\theta}\in F$ that minimize the overall loss of the data as follows:
@@ 3.2 Our proposed system - EHBS
The Embedded Hyperspectral Band Selection (EHBS) system, our proposed approach, is an end-to-end embedded system. It comprises a downstream task model enhanced with an additional layer inserted between the input layer and the task model. This added layer, is based on the principles of the Stochastic Gates (STG) model - see 2.2. Our novel adaptation is specifically tailored for hyperspectral band selection within the context of image semantic segmentation and is further detailed in 3.2.1. By adding this layer, EHBS leverages the intrinsic characteristics of hyperspectral data and addresses the unique requirements of semantic segmentation tasks, seamlessly integrating with downstream models without the need for band-selection prepossessing.
@@ 3.2.1 Adapting STG for Feature Groups and Convolutional Layers
In the setting of band selection, and in contrast to the standard feature selection setting, all features of a given band should either be included or excluded from the input. We have thus adapted the stochastic gates framework, originally designed for feature selection, to work over groups of features. This is done be altering the gates layer to either mask all of the features in the group (i.e. band) or to leave the features intact. This layer comes right after the input layer and precedes the first layer of the deep learning network of the downstream task. The Gate is applied to each band-specific 2D input in the corresponding full HSI 3D input. The output of the gate is a corresponding mask on the input based on each specific gate value.
The stochastic gates layer includes $N$ gates per each band ($N$ is the size of the spectral dimension). Each gate $s$ includes a learned parameter $\mu$ which corresponds to the significance of its corresponding input band (See Figure Figure 2). The gates layer acts as a sparse layer manipulating the input layer while preserving its shape. This is done by multiplying all corresponding feature of a band by a calculated value based on the gates $\mu$ value.
@@
Where $\Phi$ is the standard Gaussian CDF and $\mu_{i}$ is the $\mu$ value of gate $s_{i}$ and $\lambda$ is a regularization factor.
3.2.2 Integrating a parameter free optimizer
@@
From our initial experimentation, it was evident that the STG-based deep learning network is very sensitive to the setting of the learning rate. Hence, one needs to experiment and set a different learning rate value depending on the system architecture, input data and target number of bands. In order to circumvent this limitation we chose DoG, a parameter-free stochastic optimizer DoG (”Distance over Gradients”) [22]. The DoG step sizes depend on simple empirical quantities (distance from the initial point and norms of gradients) and there is no “learning rate” parameter that needs to be set.
We have used this dynamic optimizer to coup with the need to balance between the band selection objective to the segmentation loss objective in the combined training process. Employing a parameter-free stochastic method such as DoG has relieved us from the necessity of tuning the learning rate parameter thus providing a consistent model applicable to all experimental settings.
@@ 4.1 Benchmark Datasets for Evaluation
In this section, we present the benchmark datasets employed to evaluate the performance of our proposed EHBS model for Image Semantic Segmentation. We focus on two common hyperspectral semantic segmentation benchmark datasets: Pavia University (PaviaU) and Salinas that represent diverse real-world scenarios.
@@ 4.1.1 PaviaU
The PaviaU hyperspectral semantic segmentation benchmark dataset [2] captures an urban scene over Pavia, northern Italy, acquired by the ROSIS sensor during a dedicated flight campaign. This publically available dataset encompasses 103 spectral bands in wave length of 430-860 nm, and the image has dimensions of 610 by 610 pixels. It is noteworthy that certain samples in the image lack information, resulting in 42,000 valid pixels for analysis. The dataset is annotated with ground-truth labels, classifying nine distinct categories, including bitumen, asphalt, tiles, trees, and more. These annotations provide a rich foundation for evaluating the semantic segmentation performance of models across a diverse set of urban materials and features. Figure 3 displays a sample band of the scene in grey-scale and as as the corresponding color-coded annotation of the ground-truth.
@@ 4.1.2 Salinas
The Salinas hyperspectral semantic segmentation benchmark dataset [3] captures an agricultural scene in the Salinas Valley, California, gathered by the 224-band AVIRIS sensor in wave length of 430-2500 nm. Renowned for its high spatial resolution (3.7-meter pixels), the dataset covers an extensive area spanning 512 lines by 217 samples, providing approximately 54,000 valid pixels for analysis. The dataset is meticulously annotated to distinguish among 16 classes of ground truths, including vegetables, bare soils, and vineyard fields. Notably, to refine the dataset for agricultural semantic segmentation tasks, 20 water absorption bands were selectively discarded from the original dataset, specifically bands [108-112], [154-167], and 224.
@@
Figure 4 displays a sample band of the scene in grey-scale and as as the corresponding color-coded annotation of the ground-truth.
@@ 4.2 Band Selection Benchmark Methods
Although not used before for band selection, we have implemented a supervised deep learning embedded method, where $\ell_{1}$ regularization was used on bands gates values, and the k gates with the highest values were the selected ones. This implementation was inspired by similar work in which regularization was embedded in DL networks for feature selection (see 2.2) and adapted to the setting of BS by applying the regularization over the bands rather than over the individual input features. The subsequent analysis and comparison against these established methods aim to provide a robust assessment of the efficacy and innovation introduced by our proposed embedded hyperspectral band selection methods.
@@ 4.3 Model Details
Our EHBS method is easily adaptable for integration into any CNN [24], ViT [13] or other deep learning model. In practice it is a neural layer between the input and the downstream task model, which consists of gates multiplying the input values with a learned factor for every patch input data proceeded by an arbitrary downstream task model. In our comprehensive experiments, we specifically focused on the state-of-the-art CNN model proposed by Hamida et el. [8], a 3D CNN specifically tailored for spectral-spatial data extraction. We chose this model for its proven performance in hyperspectral semantic segmentation tasks. The objective of our experiments was to evaluate the effectiveness of our embedded band selection approach in comparison to non-embedded methods across varied dataset sizes, sample (patch) sizes, and cross-validation scenarios. Implementation wise, we based our code on the publicly available implementation of Hamida et el. [8] available in [1]. The corresponding architecture of the CNN network is illustrated in Figure 5. The CNN network consists of several 3D convolutional layers, proceeded by a 1D convolution and eventually followed by a final class probabilities output layer. For the hyperparameters and kernel sizes, we used the default settings consistent with the original paper and its corresponding available model implementation  [8]. We used the Pytorch deep learning framework. Batch size was set to 256. The number of epochs was set to 100 for PaviaU and 150 for Salinas based on observing the converge patterns during our initial experimentation. The initial $mu$ and $sigma$ values of the stochastic gates layer were set to the recommended value of 0.5.
We utilized the DoG optimizer, eliminating the need for explicit learning rate tuning, contributing to the adaptability and efficiency of our method.
The STG layer includes a regularization component designed to minimize the number of the active gates (or more specifically their $\mu$ values). The corresponding regularization factor ($\lambda$) was used to meet the target number of bands. We run multiple experiments with various $\lambda$ values to obtain the various required target number of bands.

### v2

@@ Abstract
The selection of hyperspectral bands plays a pivotal role in remote sensing and image analysis, with the aim of identifying the most informative spectral bands while minimizing computational overhead. This paper introduces a pioneering approach for hyperspectral band selection that offers an embedded solution, making it well-suited for resource-constrained or real-time applications. Our proposed method, embedded hyperspectral band selection (EHBS), excels in selecting the best bands without needing prior processing, seamlessly integrating with the downstream task model. This is achieved through stochastic band gates along with an approximation of the $l0$ norm on the number of selected bands as the regularization term and the integration of a dynamic optimizer, DoG, which removes the need for the required tuning of the learning rate. We conduct experiments on two distinct semantic-segmentation hyperspectral benchmark datasets, demonstrating their superiority in terms of accuracy and ease of use compared to many common and state-of-the-art methods. Furthermore, our contributions extend beyond hyperspectral band selection. Our approach’s adaptability to other tasks, especially those involving grouped features, opens promising avenues for broader applications within the realm of deep learning, such as feature selection for feature groups.
@@ 1 Introduction
While HSI offers notable advantages, it is not without its challenges. One significant drawback lies in the substantial costs associated with the physical sensor hardware. Additionally, HSI incurs increased expenses related to storing, transferring, and analyzing the considerably larger image data generated by HSI. Furthermore, the rise of very large and deep networks for vision has further increased the complexity of models and, consequently, computing costs. Hence, there is a pressing need for algorithms and methods for band selection, i.e., identifying a subset of hyperspectral bands that retains essential information for downstream tasks. Though existing research has explored band selection, the various methods proposed typically involved an unsupervised pre-processing step that is independent of the downstream HSI task, and thus, the choice of band selection may be sub-optimal.
This paper introduces Embedded Hyperspectral Band Selection (EHBS), a plug-and-play embedded method. EHBS effectively selects the optimal bands without requiring preliminary processing, seamlessly integrating with the downstream task model. EHBS utilizes an existing feature selection algorithm based on stochastic gates that was adopted to the setting of band selection of hyperspectral data. In this study, we showcase the capability of EHBS to dynamically learn optimal bands seamlessly as an integral part of the Convolutional Neural Network (CNN) implementation. This is done in the context of semantic segmentation, a visual task that predicts semantic categories for each pixel in an input image that has received growing interest in the context of hyperspectral imagery, particularly with the application of deep learning methods. Our results demonstrate that our method outperforms existing band-selection methods on two different hyperspectral semantic segmentation datasets achieving similar accuracy values of the full hyperspectral data by using only around 25% of the bands. Importantly, our approach stands out by being easily applicable to deep-learning tasks over HSI datasets.
@@ 2.1 Hyperspectral Imaging
Hyperspectral imaging captures rich spectral data per pixel, unlike standard imaging, with only three spectral samples (Red, Green, and Blue). Contemporary hyperspectral systems offer hundreds of spectral bands spanning both visible and invisible spectra, effectively creating a three-dimensional cube composed of two-dimensional grayscale images. This detailed HSI data provides insights beyond the capabilities of regular RGB imaging that can be exploited by deep learning models.
@@
Semantic segmentation is a fundamental HSI task to predict the semantic categories for each pixel of a hyperspectral input image [25]. Semantic segmentation tasks based on aerial and satellite images play an important role in a wide range of applications [20]. In recent years, the successful application of deep learning (DL) in the field of computer vision (CV) has led to a surge of work applying DL methods for data semantic segmentation, resulting in notable achievements and significant advancements.
@@
Deep learning applications for HSI have become popular and prevailed in recent years. For example, Zhang et. al [37] used a 3D CNN with transfer learning for aero image segmentation while [29] explored unsupervised hyperspectral unmixing using autoencoders to classify hyperspectral images into 3 labels (Tree, Water, Rock).
@@ 2.2 Feature Selection
Feature selection (FS) methods identify the essential features needed by a machine learning system to perform a downstream task and can be categorized broadly as filter, wrapper, and embedded methods. Filter methods eliminate irrelevant features before model learning, using statistical relevance scores [6, 10]. Wrapper methods determine feature relevance based on model performance [23, 4], but their drawback lies in computational expense [30]. Embedded methods address this by simultaneously learning the model and selecting relevant features within a self-contained and single-train organism. For example, the widely recognized Least Absolute Shrinkage and Selection Operator (LASSO) [34] is an embedded FS algorithm that minimizes loss with an $l_{1}$ constraint but is limited to linear functions. Attempts to extend LASSO using neural networks face challenges with suboptimal gradient descent on $l_{1}$ regularized objectives [15].
STG [36] proposes an embedded feature selection neural network scheme. The STG procedure is based on probabilistic relaxation of the $\ell_{0}$ norm of features or the count of the number of selected features. The $\ell_{0}$-based regularization relies on a continuous relaxation of the Bernoulli distribution; such relaxation allows the STG model to learn the parameters of the approximate Bernoulli distributions via gradient descent. The STG framework simultaneously learns either a nonlinear regression or classification function while selecting a small subset of features. It provides an information-theoretic justification for incorporating Bernoulli distribution into feature selection. Figure 1 illustrates the STG model in which the stochastic gates are attached to the $x_{i}$ input features, where the trainable parameter $\mu_{i}$ and a noise component $\epsilon_{i}$ control the choice of the gate being active or not ($z_{i}$).
@@ 2.3 Hyperspectral Band Selection
Band selection (BS) methods are categorized into two main groups. Unsupervised methods operate without using any annotations, relying on information criteria for internal information and dissimilarity from other bands. While unsupervised methods are suitable for unlabeled datasets, their lack of consideration for specific hyperspectral imaging (HSI) task performance may result in sub-optimal band choices. On the other hand, Supervised methods utilize annotated data in the selection process, tailoring band selection to the specific task and considering prediction model performance. This category is further divided into before-train (finding the best bands before training the downstream model) and embedded-in-train (training the band selector as part of the downstream model, as in our method). Following, we will describe six different methods that can be classified into five different families: Search, Ranking, Clustering, Sparsity, and DL-based. These methods were implemented and used as baseline models to be compared to our proposed model.
@@
Clustering-based methods in band selection involve grouping bands and selecting one representative from each group to minimize redundancy. Often used in conjunction with ranking, these methods aim to choose the highest-ranked representatives from clustered bands [35]. WaLuDI and WaLuMI [27] employ a distance metric based on information theory measures to assess the similarity between bands. Their approach minimizes in-cluster variance while maximizing between-cluster variance. The dissimilarity measure calculates the distance between bands using probability functions based on gray-scale pixel values. For instance, the KL divergence is employed to quantify differences between probability distributions, enhancing the effectiveness of band clustering and selection strategies.
@@
Another more recent line of work includes DL-based methods aimed at modeling the nonlinear interdependencies between the various spectral bands. BS-Nets [9] is an embedded unsupervised method that applies a DNN to reconstruct the full HSI image from partially available bands. The network’s attention mechanism was used from both spatial and spectral views to infer the best band combination for reconstruction.
@@ 3.1 Problem statement
The goal of embedded band selection methods is to simultaneously select an Indicator vector $I$ and a model $f_{\theta}\in F$ that minimize the overall loss of the data as follows:
@@ 3.2 Our proposed system - EHBS
The Embedded Hyperspectral Band Selection (EHBS) system, our proposed approach, is an end-to-end embedded system. It comprises a downstream task model enhanced with an additional layer inserted between the input layer and the task model. This added layer is based on the principles of the Stochastic Gates (STG) model - see 2.2. Our novel adaptation is specifically tailored for hyperspectral band selection within the context of image semantic segmentation and is further detailed in 3.2.1. By adding this layer, EHBS leverages the intrinsic characteristics of hyperspectral data and addresses the unique requirements of semantic segmentation tasks, seamlessly integrating with downstream models without the need for band-selection prepossessing.
@@ 3.2.1 Adapting STG for Feature Groups and Convolutional Layers
In the band selection setting, and in contrast to the standard feature selection setting, all features of a given band should either be included or excluded from the input. We have thus adapted the framework of the stochastic gate, originally designed for feature selection, to work over groups of features. This is done by altering the gates layer to either mask all of the features in the group (i.e. band) or leave the features intact. This layer follows the input layer and precedes the first layer of the deep learning network of the downstream task. The Gate is applied to each band-specific 2D input in the corresponding full HSI 3D input. The output of the gate is a corresponding mask on the input based on each specific gate value.
The stochastic gates layer includes $N$ gates per band ($N$ is the size of the spectral dimension). Each gate $s$ includes a learned parameter $\mu$, which corresponds to the significance of its corresponding input band (See Figure Figure 2). The gates layer acts as a sparse layer manipulating the input layer while preserving its shape. This is done by multiplying all corresponding features of a band by a calculated value based on the gates $\mu$ value.
@@
Where $\Phi$ is the standard Gaussian CDF, $\mu_{i}$ is the $\mu$ value of gate $s_{i}$, and $\lambda$ is a regularization factor.
3.2.2 Integrating a parameter-free optimizer
@@
From our initial experimentation, it was evident that the STG-based deep learning network is very sensitive to the setting of the learning rate. Hence, one needs to experiment and set a different learning rate value depending on the system architecture, input data, and target number of bands. In order to circumvent this limitation, we chose DoG, a parameter-free stochastic optimizer DoG (”Distance over Gradients”) [22]. The DoG step sizes depend on simple empirical quantities (distance from the initial point and norms of gradients), and there is no “learning rate” parameter that needs to be set.
We have used this dynamic optimizer to balance the band selection objective and the segmentation loss objective in the combined training process. Employing a parameter-free stochastic method such as DoG has relieved us from the necessity of tuning the learning rate parameter, thus providing a consistent model applicable to all experimental settings.
@@ 4.1 Benchmark Datasets for Evaluation
In this section, we present the benchmark datasets employed to evaluate the performance of our proposed EHBS model for Image Semantic Segmentation. We focus on two common hyperspectral semantic segmentation benchmark datasets, Pavia University (PaviaU) and Salinas, that represent diverse real-world scenarios.
@@ 4.1.1 PaviaU
The PaviaU hyperspectral semantic segmentation benchmark dataset [2] captures an urban scene over Pavia, northern Italy, acquired by the ROSIS sensor during a dedicated flight campaign. This publically available dataset encompasses 103 spectral bands in wavelengths of 430-860 nm, and the image has dimensions of 610 by 610 pixels. It is noteworthy that certain samples in the image lack information, resulting in 42,000 valid pixels for analysis. The dataset is annotated with ground-truth labels, classifying nine distinct categories, including bitumen, asphalt, tiles, trees, and more. These annotations provide a rich foundation for evaluating the semantic segmentation performance of models across a diverse set of urban materials and features. Figure 3 displays a sample band of the scene in grey scale and as the corresponding color-coded annotation of the ground truth.
@@ 4.1.2 Salinas
The Salinas hyperspectral semantic segmentation benchmark dataset [3] captures an agricultural scene in the Salinas Valley, California, gathered by the 224-band AVIRIS sensor in a wavelength of 430-2500 nm. Renowned for its high spatial resolution (3.7-meter pixels), the dataset covers an extensive area spanning 512 lines by 217 samples, providing approximately 54,000 valid pixels for analysis. The dataset is meticulously annotated to distinguish among 16 classes of ground truths, including vegetables, bare soils, and vineyard fields. Notably, to refine the dataset for agricultural semantic segmentation tasks, 20 water absorption bands were selectively discarded from the original dataset, specifically bands [108-112], [154-167], and 224.
@@
Figure 4 displays a sample band of the scene in grey scale and as the corresponding color-coded annotation of the ground truth.
@@ 4.2 Band Selection Benchmark Methods
Although not used before for band selection, we have implemented a supervised deep learning embedded method, where $\ell_{1}$ regularization was used on band gate values, and the k gates with the highest values were the selected ones. This implementation was inspired by similar work in which regularization was embedded in DL networks for feature selection (see 2.2) and adapted to the setting of BS by applying the regularization over the bands rather than over the individual input features. The subsequent analysis and comparison against these established methods aim to provide a robust assessment of the efficacy and innovation introduced by our proposed embedded hyperspectral band selection methods.
@@ 4.3 Model Details
Our EHBS method is easily adaptable for integration into any CNN [24], ViT [13] or other deep learning model. In practice, it is a neural layer between the input and the downstream task model, which consists of gates multiplying the input values with a learned factor for every patch input data proceeded by an arbitrary downstream task model. In our comprehensive experiments, we specifically focused on the state-of-the-art CNN model proposed by Hamida et el. [8], a 3D CNN specifically tailored for spectral-spatial data extraction. We chose this model for its proven performance in hyperspectral semantic segmentation tasks. The objective of our experiments was to evaluate the effectiveness of our embedded band selection approach in comparison to non-embedded methods across varied dataset sizes, sample (patch) sizes, and cross-validation scenarios. Implementation-wise, we based our code on the publicly available implementation of Hamida et al . [8] available in [1]. The corresponding architecture of the CNN network is illustrated in Figure 5. The CNN network consists of several 3D convolutional layers, preceded by a 1D convolution and eventually followed by a final class probabilities output layer. For the hyperparameters and kernel sizes, we used the default settings consistent with the original paper and its corresponding available model implementation  [8]. We used the Pytorch deep learning framework. The batch size was set to 256. The number of epochs was set to 100 for PaviaU and 150 for Salinas based on observing the converge patterns during our initial experimentation. The initial $mu$ and $sigma$ values of the stochastic gates layer were set to the recommended value of 0.5.
We utilized the DoG optimizer, eliminating the need for explicit learning rate tuning, which contributed to our method’s adaptability and efficiency.
The STG layer includes a regularization component designed to minimize the number of active gates (or, more specifically, their $\mu$ values). The corresponding regularization factor ($\lambda$) was used to meet the target number of bands. We run multiple experiments with various $\lambda$ values to obtain the various required target number of bands.